OpenStudy (anonymous):

\[\frac{ 7i-8 }{ i+1 }\] Express in simplest form, in terms of i

OpenStudy (anonymous):

I have an answer but not forsure if it is right

OpenStudy (anonymous):

(7i-8)/(i+1)=(7i-8)(i-1)/(i-1)(I-1)=(7i^2-7i-8i+1)/(i^2-1) =(7i^2-15i+1)/(-1-1)=[7(-1)-15i+1]/-2=(-6-15i)/-2 =-6/(-2)-15i/(-2)=3+15/2 i

OpenStudy (anonymous):

so it equals 18/2i